OpenStudy (anonymous):

The rectangles for a quilt design are 4.5 in by 2.5 in. What is the amount of fabric required to make eight rectangles for a quilt?

OpenStudy (anonymous):

4.5*2.5=11.25 * 8=90"^2

OpenStudy (anonymous):

thanx